OpenStudy (majorwhipage1):

Point g is located at (3,-1) and point h is located at (-2,3) find y for the point that is 2/3 the distance from point g to point h

OpenStudy (sshayer):

\[1-\frac{ 2 }{ 3 }=\frac{ 3-2 }{ 3 }=\frac{ 1 }{ 3 }\] \[\frac{ 2 }{ 3 }:\frac{ 1 }{ 3 }=2:1\] 

OpenStudy (mathstudent55):